A 25.0 mL sample of 0.150 M formic acid is titrated with a 0.150 M NaOH solution. What is the pH at the equivalence point? The Ka of formic acid is 1.8 ⋅ 10-4.
Answer #1

1 mol of NaOH reacts with 1mol of formic acid

Molarity of Formic acid (M1) = 0.150 M

Volume of formic acid (V1) = 25.0 mL

Molarity of NaOH (M2) = 0.150 M

Let the volume of NaOH required at equivalence point be V2

From Molarity equation

M1V1 = M2V2

V2 = 0.150 M * 25.0 mL / 0.150 M = 25 mL

Total volume of solution = 25 + 25 = 50.0 mL = 0.075 L

[HCOO-] = Molarity *volume = 0.150M * 0.025 L/0.050L = 0.075 moles

Kb = Kw/Ka = (10-14 ) /(1.8 *10-4) = 5.56 *10-11

HCOO- + H2O <--> HCOOH + OH-

Kb = [HCOOH][OH-] / [HCOO-]

5.56 *10-11 = x2 / (0.075)

x = 2.04 *10-6 M = [OH-]

pOH = -log [OH-] = -log (2.04 *10-6 M) = 5.69

pH = 14 - pOH = 14 -5.69 = 8.31